OpenStudy (mikezack123):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −1/2x + 3 y = −1/2x + 6 y = 2x + 3

OpenStudy (mikezack123):

@satellite73

OpenStudy (mikezack123):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

the slope of the line y = 2x+2 is m = 2 = 2/1 the perpendicular slope is found by flipping the fraction to go from 2/1 to 1/2 then flipping the sign to go from 1/2 to -1/2 This means the perpendicular slope is -1/2

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (mikezack123):

fer sure

jimthompson5910 (jim_thompson5910):

so we know that m = -1/2

jimthompson5910 (jim_thompson5910):

the line we want must pass through (6,3) which means that when x = 6, y is y = 3

jimthompson5910 (jim_thompson5910):

y = mx + b y = (-1/2)x + b ... plug in the slope we just found 3 = (-1/2)(6) + b .. plug in x = 6 and y = 3 now solve for b

jimthompson5910 (jim_thompson5910):

I'll let you take over from here

OpenStudy (mikezack123):

got it, its C

jimthompson5910 (jim_thompson5910):

Correct, it is C

OpenStudy (mikezack123):

thanks :)

jimthompson5910 (jim_thompson5910):

no problem